OpenStudy (anonymous):

what is the slope of the line (5,3) (3,1)

OpenStudy (karatechopper):

y-y/x-x is the formula for slope

OpenStudy (anonymous):

2

OpenStudy (karatechopper):

luis is right

OpenStudy (anonymous):

how when it is 2

OpenStudy (anonymous):

its -3/2

OpenStudy (karatechopper):

no no no

OpenStudy (karatechopper):

y-y/x-x=3-1/5-3 3-1=2 5-3=2 2/2 =1

OpenStudy (anonymous):

okay

OpenStudy (karatechopper):

yep

OpenStudy (karatechopper):

please click good answer..:) i will greatly appreciate it:)

OpenStudy (anonymous):

Slope is = 1

OpenStudy (anonymous):

ok well (5,3) (6,5)